ACCURATE AB INITIO QUARTIC FORCE FIELDS FOR HO_{2}^{+}

\underline{XINCHUAN HUANG} and TIMOTHY J. LEE, MS 245-6, NASA Ames Research Center, Moffett Field, California 94035..

Purely ab initio CCSD(T) quartic force fields have been determined for the lowest triplet and singlet states of HO_{2}^{+} ion with TZ-,QZ- and 5Z-level one-particle basis sets. After extrapolating to complete basis-set limit, small correction terms have now been incorporated, including core-correlation, scalar relativistic, and others. Anharmonic vibrational fundamentals and ro-vibrational spectroscopic constants are computed with perturbational and variational methods. Our best computed force field should give fundamentals to within ±5 cm^{−1}. In this talk we will present the procedure details and compare our results with previous work on this ion. Effects of small corrections and benchmark results on H_{2}O will also be discussed.